(3.3.15) Based on sample​ data, newborn males have weights with a mean of 3243.4 g and a standard deviation of 574.9 g. Newborn females have weights with a mean of 3025.6 g and a standard deviation of 724.6 g. Who has the weight that is more extreme relative to the group from which they​ came: a male who weighs 1700 g or a female who weighs 1700 ​g?

Since the z score for the male is z=? and the z score for the female is z=?​, the female/male has the weight that is more extreme.
1
Expert's answer
2020-11-25T15:05:30-0500

Given for male:

"\overline {x}= 3243.4\\s=574.9"

given for female:

"\overline{x}= 3025.6\\

s=724.6"

formula to find z score:


"Z= \frac{x- \overline{x}}{s}"

Male: Z= "\frac{1700-3243.4}{574.9}=-2.685"


Female Z="\frac{1700-3025.6}{724.6}" = -1.829


The most extreme value is the z score that is the furthest from 0. since -2.685 is more furthest than -1.829, the male who weighs 1700 is more extreme relatively.